OpenStudy (anonymous):

is (-4)1/2 a real function?

OpenStudy (swissgirl):

hmmm this is isnt a function since you r missing the x and y's its just a plain value

OpenStudy (anonymous):

and it's not a real value as it's complex

OpenStudy (anonymous):

\[\sqrt{-4}\]

OpenStudy (swissgirl):

ohh i thought it was \( 4* {1 \over 2 } \)

OpenStudy (swissgirl):

whoops \( -4* \large {1 \over 2} \)

OpenStudy (swissgirl):

@Susansamir is this the exact question?

OpenStudy (swissgirl):

Wait how did my responses go out of order

OpenStudy (anonymous):

thanks guys i got it

OpenStudy (swissgirl):

and now that i refreshed its back in order soo strange

OpenStudy (anonymous):

yes, an exponential function, (-4)1/2=-2

OpenStudy (unklerhaukus):

\[f(z)=(-4)^{1/2}\]\[\qquad=\sqrt{-4}\]\[\qquad=2i\] the function \(f(z)\) has no real component